What is the first derivative of j(x)=e* Inv x2 - 4

calculus 1/business calculus methods only
Answer #1

-X (* 1n (**) = nin x) Given that (x)= e*in(4x2-4)= e*19(x2+4)* = ze in (32-4) * In (x2-4) e* ln (x2 - 4)) (In (x2 – 4)) e*)x